OpenStudy (anonymous):

Which set of line segments with the given lengths can be formed into a right triangle? a) 24, 30, 35 b) 12, 18, 30 c) 18, 24, 30 d) 18, 24, 35

OpenStudy (cwrw238):

apply pythagoras

OpenStudy (anonymous):

i did, still couldnt find it

OpenStudy (cwrw238):

try (b again

OpenStudy (cwrw238):

30^2 = 900 24^2 = 576 18^2 = 324

hero (hero):

Hint divide row c by 3 to get 3: 4: 5

OpenStudy (cwrw238):

sorry thats c